# Efficient Inference for Large Language Model-based Generative Recommendation

### v2

@@ 4.3 In-depth Analysis
From the results as shown in Figure 3, we have the following observations. 1) the performance decline of “w/o DR” indicates that aggregating probability around the top-$K$ tokens can effectively strengthen the alignment. 2) The use of top-1 sequence from the target LLM to execute AtSpeed-R results in a lower acceptance rate. This makes sense since the top-1 sequence lacks diversity, which might overemphasize the alignment on the top-1 sequence and hurt the speedup. 3) By discarding either the constrained alignment or tree attention would lead to performance decline for both AtSpeed-S and AtSpeed-R. This is expected since removing constrained alignment would introduce potential noisy tokens, which have high probability yet will be rejected to ensure in-corpus item recommendation. Furthermore, 4) the removal of tree-based attention would hurt the speedup although it does not affect the acceptance rate, which validates the efficiency of utilizing tree-based attention to execute parallel verification.
Figure 3: Ablation study of AtSpeed-S and AtSpeed-R on Beauty dataset. Figure(a) presents the performance of removing each component of AtSpeed-S separately under strict top-$K$ verification; while figure(b) shows the results of discarding each component of AtSpeed-R separately under relaxed sampling verification.
Recommendation Accuracy of AtSpeed-R (RQ3). To examine how our relaxed verification strategy affects the recommendation accuracy, we run AtSpeed-R and the target LLM (i.e., LC-Rec (Zheng et al., 2024) based on LLaMA-7B) via sampling-based beam search. The results are presented in Table 2, from which we can observe that 1) AtSpeed-R consistently yields comparable performance compared to the target LLM when the sampling is enabled for beam search. This empirically validates the approximately equivalent output distribution of the LLM inference with and without SD. In addition, 2) while AtSpeed-R boosts the LLM inference efficiency with an over 2$\times$ speedup, it can achieve comparable or even better accuracy compared to the original top-$K$ recommendation of LC-Rec without sampling (temperature=0).
Table 2: Recommendation performance of AtSpeed-R under relaxed sampling verification on Beauty. “WS@K” denotes “walltime speedup@K”.
Temperature=0 Temperature=0.5 Temperature=1
Metrics Target LLM Target LLM AtSpeed-R Target LLM AtSpeed-R
Recall@5 0.0056 0.0057 ↑0.0001 0.0051 ↓0.0006 0.0053 ↓0.0003 0.0058 ↑0.0002
Recall@10 0.0098 0.0095 ↓0.0003 0.0091 ↓0.0007 0.0088 ↓0.0010 0.0092 ↓0.0006
WS@5 1$\times$ 1$\times$ 1.83$\times$ ↑0.83 1$\times$ 2.13$\times$ ↑1.13
WS@10 1$\times$ 1$\times$ 2.04$\times$ ↑1.04 1$\times$ 2.10$\times$ ↑1.10
@@
Figure 4: Effect of draft beam size $N$ and the alignment strength $\alpha$ on Beauty.
@@
Effect of Draft Beam Size $\bm{N}$ (RQ4). To study how the draft beam size $N$ affects the acceleration of AtSpeed, we vary $N$ from 20 to 60 and test AtSpeed-S under strict top-$K$ verification (Figure 4) and AtSpeed-R under relaxed sampling verification, respectively. From the figures, we can observe that 1) from $N=20$ to $N=60$, the acceptance rate constantly increases for both AtSpeed-S and AtSpeed-R. This is reasonable as a larger beam size can produce more drafted sequences for verification, thus enhancing the acceptance rate. Moreover, 2) the walltime speedup continues to improve steadily as $N$ gets larger. We attribute this phenomenon to the utilization of tree-based attention, which efficiently verifies drafted sequences with a single flattened sequence (refer to Appendix A.3 for detailed speedup analysis of tree-based attention). Nevertheless, 3) the improvement of acceptance rate and speedup gradually slows down (i.e., decreasing slope), indicating continuously increasing $N$ is approaching a balance between the enhanced acceptance rate and the increased overhead.
@@
Effect of Alignment Strength $\bm{\alpha}$ (RQ4). We train the draft model with different alignment strength $\alpha$ and present the results of AtSpeed-S in Figure 4(c) [^7] More analysis of AtSpeed is presented in Appendix A.3 to save space.. From the figure, it is observed that the integration of alignment loss (incrementally increased from 0) improves performance, which validates the necessity of alignment to improve acceptance rate. However, 2) blindly emphasizing the alignment loss may be ineffective, as the model struggles to capture the output of the target LLM due to the limited expressiveness of the small-sized model without the knowledge of recommendation task. Therefore, it is essential to strike a balance between the task-specific objective and the alignment objective. Empirically, we recommend setting $\alpha=0.5$ for alignment training.
@@ 5 Related Work
Speculative Decoding. As an efficient LLM decoding technique, SD (Leviathan et al., 2023; Fu et al., 2024; Du et al., 2024; Liu et al., 2024; Chen et al., 2024; Sun et al., 2024) has been proposed to reduce the LLM calls by efficiently drafting multiple future tokens and then verifying them in parallel. Early studies focus on SD in greedy decoding, where the draft model generates one token at each generation step for LLM to verify if they match the LLM’s top-$1$ token. Later on, some studies extend SD to nucleus sampling to accelerate LLM decoding without changing its output distribution (Leviathan et al., 2023). However, the early attempts simply draft one token at each step for verification, which significantly limits the acceptance rate (Xia et al., 2023). To improve the acceptance rate, subsequent work designs various relaxed verification strategies for greedy decoding such as enlarging LLM candidates with a tolerable score away from the top-$1$ token (Xia et al., 2023) and measuring the distance of the probability between the draft model and the target LLM (Kim et al., 2024). Concurrently, some studies propose different methods to draft multiple candidates for different sampling, as seen in Medusa (Cai et al., 2024), EAGLE (Li et al., 2024b) and so on. While previous work primarily applies SD for NLP tasks and follows the $N$-to-$1$ verification, our work extends SD to LLM-based generative recommendation tasks, addressing the unique challenge of $N$-to-$K$ verification, i.e., perfect matching over the top-$K$ sequences. Closely related to our work, DistillSpec (Zhou et al., 2024) investigates the use of various KD techniques to enhance the alignment between the draft model and the target LLM, to achieve high acceptance rate. However, they mainly focus on how KD helps for the $N$-to-$1$ verification. In contrast, this work studies the SD under $N$-to-$K$ verification for recommendation tasks, and proposes a framework to align the draft model with the target LLM on the top-$K$ sequences specifically for recommendation tasks.
@@
Inference Acceleration for LLM-based Recommendation. While LLM-based generative recommendations have shown remarkable performance (Li et al., 2024a), their practical implementation is impeded due to the high time latency of LLM decoding (Li et al., 2023). Consequently, accelerating LLM decoding is imperative to facilitate real-world deployment of the powerful LLM-based recommender models. To tackle this issue, several approaches leverage KD to transfer knowledge from a teacher LLM to either a smaller language model (Xu et al., 2024; Wang et al., 2024b) or a traditional recommender model (Cui et al., 2024). Additionally, some methods focus on designing more lightweight model architectures, such as narrow and deep transformer structures (Mei & Zhang, 2023). However, the application of SD is underexplored in LLM-based recommendation. DARE (Xi et al., 2024) is a relevant study that applies SD in user feature augmentation for LLM-based recommendation. However, they focus on the LLM decoding for user feature generation that requires only one response from the target LLM, which is essentially SD under the $N$-to-$1$ verification. In contrast, our work explores decoding acceleration for top-$K$ ranked item recommendation, tackling the more difficult $N$-to-$K$ verification problem. In addition to these approaches, various other techniques, such as model quantization (Xiao et al., 2023), model pruning (Hassibi & Stork, 1992), and Key-Value (KV) cache optimization (Zhang et al., 2024), can also enhance inference acceleration. Nonetheless, these methods are orthogonal to SD and can be effectively integrated with SD to further improve decoding efficiency.
@@ 6 Conclusion and Future Work
All the results in this work are reproducible. We provide all the necessary code to replicate our results in an anonymous GitHub repository. The repository includes environment configurations, run scripts, and other relevant materials. We discuss the experimental settings in Section 4.1, including implementation details such as the utilization of parameter-efficient training. Additionally, detailed experimental settings are provided in Appendix A.3.
@@ A.1 Algorithm of Relaxed Sampling Verification
A.2 Proof and detailed derivatives
With-replacement Sampling Approximation. During each step of the beam search, given the sequence probability distribution $p$ and the beam size $n$, we can effectively approximate the sampling without replacement to the sampling with replacement when the population size $N$ i.e., vocabulary size, is much larger than the sample size $n$, i.e., beam size.
First, we model the sampling with replacement with multinomial distribution, where $p_{i}$ is the probability of sampling sequence $i$. We have the mass function as
$f(X_{1}=x_{1},X_{2}=x_{2},\dots,X_{n}=x_{n};n,p_{1},p_{2},\dots,p_{n})$ (10)
$=P(X_{1}=x_{1},\dots,X_{n}=x_{n})$
$=\binom{n}{n_{1}n_{2}\dots}\prod_{i}p_{i}^{n_{i}}.$
We then model the sampling without replacement with the multivariate hypergeometric distribution. Let the population size be $N$, the subpopulation size be $N_{i}$, the sample size be $n$, and the subpopulation sampled number be $n_{i}$. Now we assume that $N\to\infty$ such that $\sum_{i}N_{i}=N$, $\sum_{i}n_{i}=n$, $\frac{N_{i}}{N}\rightarrow p_{i},\sum_{i}p_{i}=1$ as $N\to\infty$. We can write the mass function of multivariate hypergeometric distribution as
$\frac{\prod_{i}\binom{N_{i}}{n_{i}}}{\binom{N}{n}}.$ (11)
Based on the Stirling’s approximation (Marsaglia & Marsaglia, 1990), i.e., $n!\sim\sqrt{2\pi n}\binom{n}{e}^{n}$, we can rearrange Eq.(11) into
$\frac{\prod_{i}\frac{N_{i}!}{n_{i}!(N_{i}-n_{i})!}}{\frac{N!}{n!(N-n)!}}=\binom{n}{n_{1}n_{2}\dots}\prod_{i}\frac{\sqrt{N_{i}}N_{i}^{N_{i}}}{\sqrt{N}}N^{N}\prod_{i}\frac{\sqrt{N-n}(N-n)^{N-n}}{\sqrt{N_{i}-n_{i}}(N_{i}-n_{i})^{N_{i}-n_{i}}}.$ (12)
By taking the limits, $\frac{N_{i}}{N_{i}-n_{i}}\to 1$ and $\frac{N}{N-n}\to 1$, Eq.(12) is further rearranged to get the mass function for the multinomial distribution
$\binom{n}{n_{1}n_{2}\dots}\prod_{i}(\frac{N_{i}}{N})^{N_{i}}\prod_{i}(\frac{N-n}{N_{i}-n_{i}})^{(N_{i}-n_{i})}$ (13)
$=\binom{n}{n_{1}n_{2}\dots}\prod_{i}p_{i}^{n_{i}}.$
Therefore, the sampling during beam search, i.e., the sampling without replacement, can be approximated to the sampling with replacement, when the population size $N$ becomes much larger than the beam size $n$.
Derivatives of Alignment objectives of AtSpeed-S. To maximize the acceptance rate under the strict top-$K$ verification, we consider:

### v3

@@ 4.3 In-depth Analysis
From the results as shown in Figure 3, we have the following observations. 1) the performance decline of “w/o DR” indicates that aggregating probability around the top-$K$ tokens can effectively strengthen the alignment. 2) The use of top-1 sequence from the target LLM to execute AtSpeed-R results in a lower acceptance rate. This makes sense since the top-1 sequence lacks diversity, which might overemphasize the alignment on the top-1 sequence and hurt the speedup. 3) By discarding either the constrained alignment or tree attention would lead to performance decline for both AtSpeed-S and AtSpeed-R. This is expected since removing constrained alignment would introduce potential noisy tokens, which have high probability yet will be rejected to ensure in-corpus item recommendation. Furthermore, 4) the removal of tree-based attention would hurt the speedup although it does not affect the acceptance rate, which validates the efficiency of utilizing tree-based attention for verification.
Figure 3: Ablation study of AtSpeed-S and AtSpeed-R on Beauty dataset.
@@
Figure 4: Effect of draft beam size $N$ and the alignment strength $\alpha$ on Beauty dataset.
@@
Effect of Draft Beam Size $\bm{N}$ (RQ3). To study how the draft beam size $N$ affects the acceleration of AtSpeed, we vary $N$ from 20 to 60 and test AtSpeed-S under strict top-$K$ verification (Figure 4) and AtSpeed-R under relaxed sampling verification, respectively. From the figures, we can observe that 1) from $N=20$ to $N=60$, the acceptance rate constantly increases for both AtSpeed-S and AtSpeed-R. This is reasonable as a larger beam size can produce more drafted sequences for verification, thus enhancing the acceptance rate. Moreover, 2) the walltime speedup continues to improve steadily as $N$ gets larger. We attribute this phenomenon to the utilization of tree-based attention, which efficiently verifies drafted sequences with a single flattened sequence (refer to Appendix A.3 for detailed speedup analysis of tree-based attention). Nevertheless, 3) the improvement of acceptance rate and speedup gradually slows down (i.e., decreasing slope), indicating continuously increasing $N$ is approaching a balance between the enhanced acceptance rate and the increased overhead.
@@
Effect of Alignment Strength $\bm{\alpha}$ (RQ3). We train the draft model with different alignment strength $\alpha$ and present the results of AtSpeed-S in Figure 4(c) [^8] More analysis of AtSpeed is presented in Appendix A.3.2 to save space.. From the figure, it is observed that the integration of alignment loss (incrementally increased from 0) improves performance, which validates the necessity of alignment to improve acceptance rate. However, 2) blindly emphasizing the alignment loss may be ineffective, as the model struggles to capture the output of the target LLM due to the limited expressiveness of the small-sized model without the knowledge of recommendation task. Therefore, it is essential to strike a balance between the task-specific objective and the alignment objective. Empirically, we recommend setting $\alpha=0.5$ for alignment training.
@@ 5 Related Work
Speculative Decoding. As an efficient LLM decoding technique, SD (Leviathan et al., 2023; Fu et al., 2024; Du et al., 2024; Liu et al., 2024c; Chen et al., 2024; Sun et al., 2024) has been proposed to reduce the LLM calls by efficiently drafting multiple future tokens and then verifying them in parallel. Early studies focus on SD in greedy decoding, where the draft model generates one token at each generation step for LLM to verify if they match the LLM’s top-$1$ token. Later on, some studies extend SD to nucleus sampling to accelerate LLM decoding without changing its output distribution (Leviathan et al., 2023). However, the early attempts simply draft one token at each step for verification, which significantly limits the acceptance rate (Xia et al., 2023). To improve the acceptance rate, subsequent work designs various relaxed verification strategies for greedy decoding such as enlarging LLM candidates with a tolerable score away from the top-$1$ token (Xia et al., 2023) and measuring the distance of the probability between the draft model and the target LLM (Kim et al., 2024). Concurrently, some studies propose different methods to draft multiple candidates for different sampling, as seen in Medusa (Cai et al., 2024), EAGLE (Li et al., 2024a) and so on. While previous work primarily applies SD for NLP tasks and follows the $N$-to-$1$ verification, our work extends SD to LLM-based generative recommendation tasks, addressing the unique challenge of $N$-to-$K$ verification, i.e., perfect matching over the top-$K$ sequences. Closely related to our work, DistillSpec (Zhou et al., 2024) investigates the use of various KD techniques to enhance the alignment between the draft model and the target LLM, to achieve high acceptance rate. However, they mainly focus on how KD helps for the $N$-to-$1$ verification. In contrast, this work studies the SD under $N$-to-$K$ verification for recommendation tasks, and proposes a framework to align the draft model with the target LLM on the top-$K$ sequences specifically for recommendation tasks. While another conceptually similar work SpecGR (Ding et al., 2024) drafts unseen items for cold-start recommendations, this work aims at drafting beam sequences at every step for inference acceleration.
@@
Inference Acceleration for LLM-based Recommendation. While LLM-based generative recommendations have shown remarkable performance (Liu et al., 2024a; Xu et al., 2024a; Zhang et al., 2024a; Lv et al., 2025; Lin et al., 2024b; Zhao et al., 2025) compared to traditional recommender models (Liu et al., 2024b; Gao et al., 2023; Li et al., 2024b; Zhang et al., 2021), their practical implementation is impeded due to the high time latency of LLM decoding. Consequently, accelerating LLM decoding is imperative to facilitate real-world deployment of the powerful LLM-based recommender models. To tackle this issue, several approaches leverage KD to transfer knowledge from a teacher LLM to either a smaller language model (Xu et al., 2024b; Wang et al., 2024b) or a traditional recommender model (Cui et al., 2024). Additionally, some methods focus on designing more lightweight model architectures, such as narrow and deep transformer structures (Mei & Zhang, 2023). However, the application of SD is underexplored in LLM-based recommendation. DARE (Xi et al., 2024) is a relevant study that applies SD in user feature augmentation for LLM-based recommendation. However, they focus on the LLM decoding for user feature generation that requires only one response from the target LLM, which is essentially SD under the $N$-to-$1$ verification. In contrast, our work explores decoding acceleration for top-$K$ ranked item recommendation, tackling the more difficult $N$-to-$K$ verification problem. In addition to these approaches, various other techniques, such as model quantization (Xiao et al., 2023), model pruning (Hassibi & Stork, 1992), and Key-Value (KV) cache optimization (Zhang et al., 2024b), can also enhance inference acceleration. Nonetheless, these methods are orthogonal to SD and can be effectively integrated with SD to further improve decoding efficiency.
@@ 6 Conclusion and Future Work
All the results in this work are reproducible. We provide all the necessary code to replicate our results in an GitHub repository. The repository includes environment configurations, run scripts, and other relevant materials. We discuss the experimental settings in Section 4.1, including implementation details such as the utilization of parameter-efficient training. Additionally, detailed experimental settings are provided in Appendix A.3.1.
@@ A.1 Algorithm of Relaxed Sampling Verification
In Algorithm 2, $\text{norm}(\max(0,p_{j}-q_{j}))=\frac{\max(0,p_{j}-q_{j})}{\sum\max(0,p_{j}-q_{j})}$.
A.2 Proof and detailed derivation
A.2.1 With-replacement sampling approximation
We aim to prove that the distribution of sampling without replacement is approximately equivalent to that of sampling with replacement. Our proof is mainly based on Stirling’s approximation (Dutka, 1991). In the following, we will first clarify notations, and introduce multivariate hypergeometric distribution and the multinomial distribution, which are used to model the sampling without replacement and sampling with replacement, respectively. We then present Stirling’s approximation, and show the step-by-step proof.
Notations. To model the sampling, we have the total population size $N$, sample size $n$, the category size $r$, the number of items in category $i$ in the population $K_{i}$, and the number of items in category $i$ in the samples $k_{i}$. In the case of sequence sampling in LLM decoding, the population includes every possible sequence. Every possible sequence is a unique category, and the sample size $n$ is the beam size, the population size is the total number of all possible sequences at each beam search step. Now we assume the population sizes go to infinity in such a way that $p_{i}=\frac{K_{i}}{N}$. And we have $\sum_{i=1}^{r}k_{i}=n$ and $\sum_{i=1}^{r}K_{i}=N$.
Multivariate Hypergeometric Distribution. Formally, when sampling without replacement, the probability of drawing $k_{1},k_{2},\dots,k_{r}$ items from each category is given by the multivariate hypergeometric distribution
$P_{\text{hyper}}(k_{1},k_{2},\dots,k_{r})$ $=\frac{\prod_{i=1}^{r}\binom{K_{i}}{k_{i}}}{\binom{N}{n}}$ (10)
$=\frac{\prod_{i=1}^{r}\frac{K_{i}!}{k_{i}!(K_{i}-k_{i})!}}{\frac{N!}{n!(N-n)!}}.$
Multinomial Distribution. Formally, when sampling with replacement, the probability follows the multinomial distribution as
$P_{\text{multi}}(k_{1},k_{2},\dots,k_{r})=\frac{n!}{\prod_{i=1}^{r}k_{i}!}\prod_{i=1}^{r}p_{i}^{k_{i}}.$ (11)
Stirling’s Approximation. Stirling’s approximation (Dutka, 1991) gives us the approximation of the logarithm of factorials as:
$\ln n!\approx n\ln n-n+\frac{1}{2}\ln(2\pi n).$ (12)
Theorem 1.
When population size $N$ is large and sample size $n$ is small compared to $N$ (i.e., $n\ll N$), the multivariate hypergeometric distribution approximates the multinomial distribution:
$P_{\text{hyper}}(k_{1},k_{2},\dots,k_{r})\approx P_{\text{multi}}(k_{1},k_{2},\dots,k_{r}).$ (13)
Proof.
We can expand the logarithm of multivariate hypergeometric probability in factorials as follows:
$\ln P_{\text{hyper}}(k_{1},k_{2},\dots,k_{r})=\ln\frac{\prod_{i=1}^{r}\frac{K_{i}!}{k_{i}!(K_{i}-k_{i})!}}{\frac{N!}{n!(N-n)!}}$ (14)
$=$ $\ln\prod_{i=1}^{r}\frac{K_{i}!}{k_{i}!(K_{i}-k_{i})!}-\ln\frac{N!}{n!(N-n)!}$
$=$ $\sum_{i=1}^{r}\ln\frac{K_{i}!}{k_{i}!(K_{i}-k_{i})!}-\ln\frac{N!}{n!(N-n)!}$
$=$ $\sum_{i=1}^{r}[\ln K_{i}!-\ln k_{i}!-\ln(K_{i}-k_{i})!]-[\ln N!-\ln n!-\ln(N-n)!].$
Using the Stirling’s approximation, we have:
[…]
Then, we can substitute the logarithm of factorials with the approximation in Eq.(14) as:
$\ln P_{\text{hyper}}(k_{1},k_{2},\dots,k_{r})$ (16)
$=$ $\sum_{i=1}^{r}[\ln K_{i}!-\ln k_{i}!-\ln(K_{i}-k_{i})!]-[\ln N!-\ln n!-\ln(N-n)!]$
$=$ $\sum_{i=1}^{r}[K_{i}\ln K_{i}-K_{i}+\frac{1}{2}\ln 2\pi K_{i}-(k_{i}\ln k_{i}-k_{i}+\frac{1}{2}\ln 2\pi k_{i})$
$\quad\quad-((K_{i}-k_{i})\ln(K_{i}-k_{i})-(K_{i}-k_{i})+\frac{1}{2}\ln 2\pi(K_{i}-k_{i}))]$
$\quad\quad-[N\ln N-N+\frac{1}{2}\ln 2\pi N-(n\ln n-n+\frac{1}{2}\ln 2\pi n)$
$\quad\quad-((N-n)\ln(N-n)-(N-n)+\frac{1}{2}\ln 2\pi(N-n))]$
$=$ $\sum_{i=1}^{r}[K_{i}\ln K_{i}-k_{i}\ln k_{i}-(K_{i}-k_{i})\ln(K_{i}-k_{i})+\frac{1}{2}\ln 2\pi K_{i}-\frac{1}{2}\ln 2\pi k_{i}-\frac{1}{2}\ln 2\pi(K_{i}-k_{i})]$
$\quad\quad-[N\ln N-n\ln n-(N-n)\ln(N-n)+\frac{1}{2}\ln 2\pi N-\frac{1}{2}\ln 2\pi n-\frac{1}{2}\ln 2\pi(N-n)].$
Since $N$ is a very large number and $n\ll N$, $k_{i}\ll K_{i}$, we have $\frac{1}{2}\ln 2\pi K_{i}-\frac{1}{2}\ln 2\pi k_{i}-\frac{1}{2}\ln 2\pi(K_{i}-k_{i})\approx 0$ and $\frac{1}{2}\ln 2\pi N-\frac{1}{2}\ln 2\pi n-\frac{1}{2}\ln 2\pi(N-n)\approx 0$. Then, the logarithm of multivariate hypergeometric distribution is approximated as:
$\ln P_{\text{hyper}}\approx\sum_{i=1}^{r}[K_{i}\ln K_{i}-k_{i}\ln k_{i}-(K_{i}-k_{i})\ln(K_{i}-k_{i})]-[N\ln N-n\ln n-(N-n)\ln(N-n)].$ (17)
Since $k_{i}$ is small compared to $K_{i}$, we can expand $\ln(K_{i}-k_{i})$ using Taylor expansion
$\ln(K_{i}-k_{i})=\ln K_{i}-\frac{k_{i}}{K_{i}}-\frac{1}{2}(\frac{k_{i}}{K_{i}})^{2}+\dots,$ (18)
where we can neglect the high-order terms and obtain
$\ln(K_{i}-k_{i})\approx\ln K_{i}-\frac{k_{i}}{K_{i}}.$ (19)
Similarly, for $\ln(N-n)$, we have
$\ln(N-n)\approx\ln N-\frac{n}{N}.$ (20)
We can then substitute $\ln(K_{i}-k_{i})$ and $\ln(N-n)$ in logarithm of multivariate hypergeometric distribution (Eq.(17)) and obtain
$\ln P_{\text{hyper}}$ $\approx\sum_{i=1}^{r}[K_{i}\ln K_{i}-k_{i}\ln k_{i}-(K_{i}-k_{i})\ln(K_{i}-k_{i})]-[N\ln N-n\ln n-(N-n)\ln(N-n)]$ (21)
$=\sum_{i=1}^{r}[K_{i}\ln K_{i}-k_{i}\ln k_{i}-(K_{i}-k_{i})(\ln K_{i}+\frac{k_{i}}{K_{i}})]-[N\ln N-n\ln n-(N-n)(\ln N+\frac{n}{N})]$
$=\sum_{i=1}^{r}[K_{i}\ln K_{i}-k_{i}\ln k_{i}-(K_{i}\ln K_{i}-k_{i}\ln K_{i}+k_{i}-\frac{k_{i}^{2}}{K_{i}})]-[N\ln N-n\ln n-(N\ln N-n\ln N+n-\frac{n^{2}}{N})]$
$=\sum_{i=1}^{r}[k_{i}\ln\frac{K_{i}}{k_{i}}+\frac{k_{i}^{2}}{K_{i}}-k_{i}]-[n\ln\frac{N}{n}+\frac{n^{2}}{N}-n]$
$=\sum_{i=1}^{r}[k_{i}\ln\frac{K_{i}}{k_{i}}+\frac{k_{i}^{2}}{K_{i}}]-[n\ln\frac{N}{n}+\frac{n^{2}}{N}].\quad\quad(\text{we have}-\sum_{i=1}^{r}k_{i}+n=0\text{ in last expression})$
We then relate $K_{i}$ to $N$ and $p_{i}$. Since $K_{i}=Np_{i}$, we have $\ln\frac{K_{i}}{k_{i}}=\ln\frac{Np_{i}}{k_{i}}$. We also have $n=\sum_{i=1}^{r}k_{i}$. Then, we can express the logarithm of hypergeometric distribution in terms of $p_{i}$ and $k_{i}$ as
$\ln P_{\text{hyper}}$ $\approx\sum_{i=1}^{r}[k_{i}\ln\frac{K_{i}}{k_{i}}+\frac{k_{i}^{2}}{K_{i}}]-[n\ln\frac{N}{n}+\frac{n^{2}}{N}]$ (22)
$=\sum_{i=1}^{r}[k_{i}\ln\frac{Np_{i}}{k_{i}}+\frac{k_{i}^{2}}{Np_{i}}]-[n\ln\frac{N}{n}+\frac{n^{2}}{N}]$
$=\sum_{i=1}^{r}[k_{i}(\ln N+\ln p_{i}-\ln k_{i})+\frac{k_{i}^{2}}{Np_{i}}]-[n\ln\frac{N}{n}+\frac{n^{2}}{N}]$
$=\sum_{i=1}^{r}k_{i}\ln N+\sum_{i=1}^{r}[k_{i}(\ln p_{i}-\ln k_{i})+\frac{k_{i}^{2}}{Np_{i}}]-n\ln\frac{N}{n}-\frac{n^{2}}{N}$
$=n\ln N-n\ln N+n\ln n-\frac{n^{2}}{N}+\sum_{i=1}^{r}[k_{i}(\ln p_{i}-\ln k_{i})+\frac{k_{i}^{2}}{Np_{i}}]\quad(\text{we have }\sum_{i=1}^{r}k_{i}=n\text{ in last expression})$
$=\sum_{i=1}^{r}[k_{i}(\ln p_{i}-\ln k_{i})+\frac{k_{i}^{2}}{Np_{i}}]+n\ln n-\frac{n^{2}}{N}$
$=n\ln n-\sum_{i=1}^{r}k_{i}\ln k_{i}+\sum_{i=1}^{r}k_{i}\ln p_{i}-\frac{n^{2}}{N}+\sum_{i=1}^{r}\frac{k_{i}^{2}}{Np_{i}}.$
Now the approximation of the logarithm of multivariate hypergeometric distribution has been finished. Similarly, we approximate the multinomial distribution with Stirling’s approximation. We expand the logarithm of multinomial distribution as
$\ln P_{\text{multi}}(k_{1},k_{2},\dots,k_{r})$ $=\ln\frac{n!}{\prod_{i=1}^{r}k_{i}!}\prod_{i=1}^{r}p_{i}^{k_{i}}$ (23)
$=\ln n!-\sum_{i=1}^{r}\ln k_{i}!+\sum_{i=1}^{r}k_{i}\ln p_{i}.$
Using Stirling’s approximation, we have $\ln n!\approx n\ln n-n$ and $\ln k_{i}!\approx k_{i}\ln k_{i}-k_{i}$. We then substitute $\ln n!$ and $\ln k_{i}!$ with the approximation and obtain
$\ln P_{\text{multi}}$ $\approx n\ln n-n-\sum_{i=1}^{r}(k_{i}\ln k_{i}-k_{i})+\sum_{i=1}^{r}k_{i}\ln p_{i}$ (24)
$=n\ln n-n-\sum_{i=1}^{r}k_{i}\ln k_{i}+\sum_{i=1}^{r}k_{i}+\sum_{i=1}^{r}k_{i}\ln p_{i}.$
Since we have $\sum_{i=1}^{r}k_{i}=n$, we have
$\ln P_{\text{multi}}$ $\approx n\ln n-n-\sum_{i=1}^{r}k_{i}\ln k_{i}+n+\sum_{i=1}^{r}k_{i}\ln p_{i}$ (25)
$=n\ln n-\sum_{i=1}^{r}k_{i}\ln k_{i}+\sum_{i=1}^{r}k_{i}\ln p_{i}.$
Now, comparing the approximated logarithm of multinomial distribution with the approximated logarithm of multivariate hypergeometric distribution, we have
$\ln P_{\text{multi}}\approx\ln P_{\text{hyper}}+\frac{n^{2}}{N}-\sum_{i=1}^{r}\frac{k_{i}^{2}}{Np_{i}}.$ (26)
Note that the term $\frac{n^{2}}{N}-\sum_{i=1}^{r}\frac{k_{i}^{2}}{Np_{i}}$ is negalectable when $N$ is large and $k_{i}$ and $n$ are small compared to N. Therefore, we show that when the population size $N$ is large (e.g., all sequences for sampling) and $k_{i},n$ are small (e.g., $k_{i}=1$ or 00 since each sequence represents a category and $n$ is usually less than 20 in LLM-based recommendation), the multivariate hypergeometric distribution can be approximated to the multinomial distribution. That is, sampling without replacement is approximately equivalent to sampling with replacement. ∎
A.2.2 Derivation of Alignment objectives of AtSpeed-S
To maximize the acceptance rate under the strict top-$K$ verification, we consider: